# From Code to Correctness: Closing the Last Mile of Code Generation with Hierarchical Debugging

### v2

@@ 3.5 Bottom-up Debugging
We initiate the process by calling MGDebugger on the main function with the decomposed code $f_{\text{root}}$ and the set of public test cases $\mathcal{T}_{\text{pub}}$. MGDebugger traverses the hierarchical structure in a depth-first manner, recursively debugging each subfunction before moving on to the higher-level functions. For each specific subfunction, MGDebugger generates relevant test cases and debugs the function based on the results. When a fix is identified, MGDebugger updates the function and propagates the changes to the dependent functions. This recursive, bottom-up strategy systematically addresses bugs , beginning with the most granular levels and progressively advancing through the function hierarchy. This method accommodates various types of bugs at different abstraction levels, from low-level syntax errors to high-level logical flaws, by focusing on one level of the hierarchy at a time and building up the corrected code in a structured manner. The detailed algorithm is presented in Algorithm 1.
4 Experiments
4.1 Setup
@@
Models We select three state-of-the-art LLMs ranging from 7B to 22B parameters as backbones for code generation and debugging: CodeQwen1.5 (7B) (Bai et al., 2023), DeepSeek-Coder-V2-Lite (16B) (Zhu et al., 2024), and Codestral (22B) (Mistral AI team, 2024). Please refer to Appendix C for our implementation details.
@@
Datasets We conduct experiments on three datasets. HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) are two widely used benchmarks for evaluating code generation systems with 164 and 500 problems, respectively. The HumanEvalFix dataset (Muennighoff et al., 2023) consists of 164 buggy functions with six different bug categories: value misuse, missing logic, excess logic, operator misuse, variable misuse, and function misuse. The detailed explanations and distribution of bug categories can be found in Appendix B.
@@
Metrics We adopt two metrics to evaluate our method: 1) Accuracy (Chen et al., 2023b; Zhong et al., 2024), which measures the overall proportion of correct code samples among all generated code samples after debugging. A code is correct iff it passes all private test cases assigned to it. 2) Repair Success Rate (RSR) (Yasunaga & Liang, 2021), which refers to the proportion of fixed code samples to the total number of buggy code samples.
@@
Baselines We compare MGDebugger with eight state-of-the-art methods for debugging LLM-generated code. 1) Simple Feedback is a basic baseline that informs the LLM that the code is incorrect and asks it to fix the issue. 2) Self-Edit (Zhang et al., 2023) prompts the LLM to edit the code based on the execution results of the test cases. 3) Self-Debugging (Chen et al., 2023b) has two variants: Self-Debugging (Expl.) prompts the LLM to explain the generated code line-by-line, while Self-Debugging (Trace) asks the LLM to dry run the code for debugging. 4) LDB (Zhong et al., 2024) segments the code into basic blocks, functions or lines, and tracks variable values during runtime after each block to verify correctness against the task description. 5) Reflexion (Shinn et al., 2023) asks the LLM to reflect on the previous code given execution results and uses a memory buffer to enable iterative refinement.
Table 1: Results of MGDebugger and other methods on HumanEval and MBPP. Acc.: Accuracy, $\Delta$: Improvement over baseline (No-Debugging), RSR: Repair Success Rate.
Model Method Dataset
@@
Acc. $\Delta$ Acc. RSR Acc. $\Delta$ Acc. RSR
DeepSeek-Coder-V2-Lite No-Debugging 76.8 67.2
Simple Feedback 82.3 +5.5 23.7 69.4 +2.2 6.7
Self-Edit 82.9 +6.1 26.3 71.2 +4.0 12.2
LDB (Block) 84.1 +7.3 31.6 74.0 +6.8 20.7
LDB (Line) 82.3 +5.5 23.7 71.8 +4.6 14.0
LDB (Function) 81.7 +4.9 21.1 72.6 +5.3 16.5
Self-Debugging (Expl.) 87.2 +10.4 44.7 73.4 +6.2 18.9
Self-Debugging (Trace) 86.0 +9.2 39.5 72.6 +5.3 16.5
Reflexion 90.9 +14.1 60.5 76.6 +9.4 28.7
MGDebugger 94.5 +17.7 76.3 80.0 +12.8 39.0
CodeQwen1.5 No-Debugging 76.2 67.4
Simple Feedback 85.4 +9.2 38.5 74.0 +6.6 20.2
Self-Edit 84.1 +7.9 33.3 75.0 +7.6 23.3
LDB (Block) 79.3 +3.1 12.8 72.8 +5.4 16.6
LDB (Line) 79.9 +3.7 15.4 72.6 +5.2 16.0
LDB (Function) 80.5 +4.3 17.9 72.8 +5.4 16.6
Self-Debugging (Expl.) 87.8 +11.6 48.7 77.4 +10.0 30.7
Self-Debugging (Trace) 84.8 +8.6 35.9 76.8 +9.4 28.8
Reflexion 87.8 +11.6 48.7 78.6 +11.2 34.4
MGDebugger 91.5 +15.3 64.1 80.8 +13.4 41.1
Codestral No-Debugging 75.6 65.4
Simple Feedback 88.4 +12.8 52.5 71.6 +6.2 17.9
Self-Edit 86.0 +10.4 42.5 75.8 +10.4 30.0
LDB (Block) 83.5 +7.9 32.5 72.2 +6.8 19.7
LDB (Line) 83.5 +7.9 32.5 71.8 +6.4 18.5
LDB (Function) 82.3 +6.7 27.5 72.0 +6.6 19.1
Self-Debugging (Expl.) 89.6 +14.0 57.5 76.4 +11.0 31.8
Self-Debugging (trace) 84.1 +8.5 35.0 73.6 +8.2 23.7
Reflexion 86.6 +11.0 45.0 75.2 +9.8 28.3
MGDebugger 94.5 +18.9 77.5 76.8 +11.4 32.9
4.2 Main Results
@@
Moreover, MGDebugger demonstrates remarkable debugging capabilities, particularly with DeepSeek-Coder-V2-Lite (16B) and Codestral (22B), where it achieves an accuracy of 94.5% on the HumanEval dataset, the highest score among all methods. This is especially impressive considering that MGDebugger operates in a zero-shot setting without task-specific retraining. This result illustrates the inherent debugging ability of larger LLMs with MGDebugger. Additionally, the method’s performance on MBPP, achieving an RSR of up to 41.1% with smaller models like CodeQwen1.5 (7B), further underscores its robustness. In general, these results validate MGDebugger as a highly effective and scalable debugging method for LLM-generated code.
4.3 Ablation Study
To understand the contribution of each component in MGDebugger and validate our design choices, we conduct an ablation study by systematically removing key components of our method: hierarchical code decomposition, LLM-simulated execution, and test case generation for subfunction debugging. Each variant is evaluated on both the HumanEval and MBPP datasets using the DeepSeek-Coder-V2-Lite model.
Table 2: Ablation study results for DeepSeek-Coder-V2-Lite. Acc.: Accuracy, $\Delta$ Acc.: Improvement over baseline (No-Debugging), RSR: Repair Success Rate.
@@ 4.3 Ablation Study
Acc. (%) $\Delta$ Acc. (%) RSR (%) Acc. (%) $\Delta$ Acc. (%) RSR (%)
MGDebugger 94.5 +17.7 76.3 80.0 +12.8 39.0
- w/o Hierarchical Debugging 89.0 +12.2 52.6 78.2 +11.0 33.5
- w/o Simulated Execution 90.2 +13.4 61.3 79.2 +12.0 36.6
- w/o Test Case Generation 90.9 +14.1 60.5 79.2 +12.0 36.6
@@
4.4 Debugging Different Types of Bugs
Table 3: Performance (RSR) on different bug categories in HumanEvalFix with different models. The best and second-best scores are highlighted in bold and underline, respectively.
@@ 4.4 Debugging Different Types of Bugs
To assess the versatility and effectiveness of MGDebugger across various bug categories, we carry out experiments using the HumanEvalFix dataset, which is specifically designed to evaluate code debugging performance. The dataset involves six distinct bug categories: value misuse, missing logic, excess logic, operator misuse, variable misuse, and function misuse, allowing us to examine how effectively MGDebugger addresses different types of programming errors compared to existing methods. The detailed explanations of each bug category are available in Appendix B.
Table 3 presents the RSRs across various bug categories. We observe that MGDebugger consistently outperforms other methods with significantly higher overall accuracies. And MGDebugger achieves a remarkable repair success rate of 97.6% using DeepSeek-Coder, with 100% success rates in all bug categories except for value misuse. This is particularly notable given the complexity and diversity of the bugs in the dataset. This highlights the effectiveness of the hierarchical debugging strategy.
@@
4.5 Debugging Code with Varying Length
Figure 3: Repair success rate of different methods when debugging code of different lengths on HumanEvalFix with DeepSeek-Coder. MGDebugger consistently outperforms other methods across different code lengths, especially in long codes.
Figure 4: Impact of the number of debugging attempts on the cumulative repair success rate of MGDebugger and other methods on HumanEvalFix with DeepSeek-Coder. MGDebugger continues to improve with more debug attempts and achieves the highest success rate.
@@ 4.5 Debugging Code with Varying Length
We further assess the versatility of MGDebugger in debugging code of varing lengths (i.e., number of tokens), since code length often correlates with complexity and debugging challenges. We categorize code snippets from the HumanEvalFix dataset into short, medium, and long groups, ensuring equal sample sizes. We subsequently analyze the RSR scores obtained by MGDebugger and baselines when using DeepSeek-Coder as the backbone LLM.
@@
The results are presented in Figure 4. We can observe that as the code length increases, most methods experience an obvious decrease in performance due to the increased complexity. We note that MGDebugger consistently outperforms other methods in different code lengths and especially excels in debugging longer and more complex code snippets. This showcases the scalability and robustness of MGDebugger in handling code of varying lengths and complexities. The results on other two datasets are available in Appendix D, where MGDebugger also consistently outperforms other methods across different code lengths.
4.6 Impact of Debug Attempts
@@
Another important factor for LLM-based debugging is the number of debugging attempts. Iterative debugging allows LLMs to refine their corrections over multiple passes, potentially leading to better outcomes. We aim to assess MGDebugger’s ability to improve over successive iterations. Following Zhong et al. (2024), we vary the number of debugging attempts from 1 to 10 using the HumanEvalFix dataset and DeepSeek-Coder.
The results in Figure 4 show that MGDebugger achieves the highest cumulative RSR score among all methods, highlighting its ability to continually refine its debugging over multiple attempts. In particular, while most methods plateau after the first few debug attempts, MGDebugger and Reflexion continue to improve with more iterations. This result underscores the great potential of MGDebugger for iterative and comprehensive debugging, making it a promising solution for complex and challenging code repair tasks. The results on the other two datasets are available in Appendix E, where MGDebugger outperforms other methods from the first attempt and continues to improve with great potential.
4.7 Case Study
Figure 5: Examples of code debugging by various methods on HumanEvalFix with DeepSeek-Coder. The three baseline methods fix the original bug but introduce new bugs that will miss the last “1” in the results. By contrast, MGDebugger successfully identifies and corrects the bug after decomposing the code into clear subfunctions for separate debugging.

### v3

@@ 2.5. Bottom-up Debugging
We initiate the process by calling MGDebugger on the main function with the decomposed code $f_{\text{root}}$ and the set of public test cases $\mathcal{T}_{\text{pub}}$. MGDebugger traverses the hierarchical structure in a depth-first manner, recursively debugging each subfunction before moving on to the higher-level functions. For each specific subfunction, MGDebugger generates relevant test cases and debugs the function based on the results. When a fix is identified, MGDebugger updates the function and propagates the changes to the dependent functions. This recursive, bottom-up strategy systematically addresses bugs by beginning with the most granular levels and progressively advancing through the function hierarchy. This method accommodates various types of bugs at different abstraction levels, from low-level syntax errors to high-level logical flaws, by focusing on one level of the hierarchy at a time and building up the corrected code in a structured manner. The detailed algorithm is presented in Algorithm 1.
3. Experiments
We conduct experiments to evaluate the effectiveness of MGDebugger, aiming to answer the following research questions:
  • RQ1: How effective is MGDebugger in debugging LLM-generated code?
  • RQ2: To what extent do individual components of MGDebugger contribute to its overall performance?
  • RQ3: How does MGDebugger perform in fixing bugs across different categories?
  • RQ4: How consistent is MGDebugger’s performance across different debugging scenarios?
  • RQ5: How will MGDebugger generalize to fixing real-world software defects?
3.1. Setup
@@
Models We select three state-of-the-art open-source LLMs ranging from 7B to 22B parameters as backbones for code generation and debugging in the experiments: CodeQwen1.5 (7B) (Team, 2023), DeepSeek-Coder-V2-Lite (16B) (AI, 2024), and Codestral (22B) (Mistral AI team, 2024). To validate the generalizability of MGDebugger, stronger general-prupose models like GPT-4o (OpenAI, 2023), Claude 3.5 Sonnet (Anthropic, 2024), and LLaMA 3.1 70B (Touvron et al., 2023) are also included in the Experiments.
@@
Datasets We conduct experiments on three widely used benchmarks for LLM-based code generation and debugging. HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) are two widely used benchmarks for evaluating code generation systems with 164 and 500 problems, respectively. The HumanEvalFix dataset (Muennighoff et al., 2023) consists of 164 buggy functions with six different bug categories: value misuse, missing logic, excess logic, operator misuse, variable misuse, and function misuse.
@@
Metrics We adopt two metrics to evaluate our method: 1) Accuracy (Chen et al., 2023b; Zhong et al., 2024), which measures the overall proportion of correct code samples among all generated code samples after debugging. A code is correct iff it passes all private test cases assigned to it. 2) Repair Success Rate (RSR) (Yasunaga and Liang, 2021), which refers to the proportion of fixed code samples to the total number of buggy code samples. The former metric indicates to which extent MGDebugger can improve the code generation systems, and the latter metric directly reflects the effectiveness of MGDebugger in fixing bugs from LLM-generated code.
@@
Baselines We compare MGDebugger with eight state-of-the-art methods for debugging LLM-generated code. 0) No-Debugging baseline refers to directly using the LLM to generate code without applying any debugging method, following (Chen et al., 2023b; Zhong et al., 2024). This baseline creates the initial set of seed programs, the subset of buggy programs then become targets for various debugging methods. 1) Simple Feedback is a basic baseline that informs the LLM that the code is incorrect and asks it to fix the issue. 2) Self-Edit (Zhang et al., 2023) prompts the LLM to edit the code based on the execution results of the test cases. 3) Self-Debugging (Chen et al., 2023b) has two variants: Self-Debugging (Expl.) prompts the LLM to explain the generated code line-by-line, while Self-Debugging (Trace) asks the LLM to dry run the code for debugging. 4) LDB (Zhong et al., 2024) segments the code into basic blocks, functions or lines, and tracks variable values during runtime after each block to verify correctness against the task description. 5) Reflexion (Shinn et al., 2023) asks the LLM to reflect on the previous code given execution results and uses a memory buffer to enable iterative refinement.
Implementation Details We generate seed programs for HumanEval and MBPP using the BigCode Evaluation Harness framework [^3] https://github.com/bigcode-project/bigcode-evaluation-harness. The specific versions of models used in our experiments are DeepSeek-Coder-V2-Lite-Instruct, CodeQwen1.5-7B-Chat, and Codestral-22B-v0.1. All experiments are conducted on NVIDIA A100 GPUs with 80GB memory. During debugging, we use the vLLM engine [^4] https://github.com/vllm-project/vllm to serve the LLMs, setting the maximum token length according to each LLM’s max length. Following Zhong et al. (2024), we limit the maximum number of debugging iterations to 10 for all methods on HumanEval, MBPP, and HumanEvalFix. To obtain visible test cases for HumanEval and HumanEvalFix, we extract the given visible test cases from the task description. For MBPP, we use the first test case of each problem as the visible test case and use the rest as hidden test cases, in line with the settings referenced from Chen et al. (2023b) and Zhong et al. (2024). Additionally, the sampling temperature that controls the randomness of LLM decoding is set to 0.8 in MGDebugger. For baseline methods that incorporate external feedback from the environments, we follow (Zhong et al., 2024) to execute the complete test code to collect runtime feedback.
3.2. Main Results (RQ1)
Table 1. Main results.
Method Dataset
@@ 3.2. Main Results (RQ1)
Acc. (%) RSR (%) Acc. (%) RSR (%)
DeepSeek-Coder-V2-Lite
No-Debugging 76.8 67.2
Simple Feedback 82.3 (+5.5) 23.7 69.4 (+2.2) 6.7
Self-Edit 82.9 (+6.1) 26.3 71.2 (+4.0) 12.2
LDB (Block) 84.1 (+7.3) 31.6 74.0 (+6.8) 20.7
LDB (Line) 82.3 (+5.5) 23.7 71.8 (+4.6) 14.0
LDB (Function) 81.7 (+4.9) 21.1 72.6 (+5.3) 16.5
Self-Debugging (Expl.) 87.2 (+10.4) 44.7 73.4 (+6.2) 18.9
Self-Debugging (Trace) 86.0 (+9.2) 39.5 72.6 (+5.3) 16.5
Reflexion 90.9 (+14.1) 60.5 76.6 (+9.4) 28.7
MGDebugger 94.5 (+17.7) 76.3 80.0 (+12.8) 39.0
CodeQwen1.5
No-Debugging 76.2 67.4
Simple Feedback 85.4 (+9.2) 38.5 74.0 (+6.6) 20.2
Self-Edit 84.1 (+7.9) 33.3 75.0 (+7.6) 23.3
LDB (Block) 79.3 (+3.1) 12.8 72.8 (+5.4) 16.6
LDB (Line) 79.9 (+3.7) 15.4 72.6 (+5.2) 16.0
LDB (Function) 80.5 (+4.3) 17.9 72.8 (+5.4) 16.6
Self-Debugging (Expl.) 87.8 (+11.6) 48.7 77.4 (+10.0) 30.7
Self-Debugging (Trace) 84.8 (+8.6) 35.9 76.8 (+9.4) 28.8
Reflexion 87.8 (+11.6) 48.7 78.6 (+11.2) 34.4
MGDebugger 91.5 (+15.3) 64.1 80.8 (+13.4) 41.1
Codestral
No-Debugging 75.6 65.4
Simple Feedback 88.4 (+12.8) 52.5 71.6 (+6.2) 17.9
Self-Edit 86.0 (+10.4) 42.5 75.8 (+10.4) 30.0
LDB (Block) 83.5 (+7.9) 32.5 72.2 (+6.8) 19.7
LDB (Line) 83.5 (+7.9) 32.5 71.8 (+6.4) 18.5
LDB (Function) 82.3 (+6.7) 27.5 72.0 (+6.6) 19.1
Self-Debugging (Expl.) 89.6 (+14.0) 57.5 76.4 (+11.0) 31.8
Self-Debugging (trace) 84.1 (+8.5) 35.0 73.6 (+8.2) 23.7
Reflexion 86.6 (+11.0) 45.0 75.2 (+9.8) 28.3
MGDebugger 94.5 (+18.9) 77.5 76.8 (+11.4) 32.9
@@
Moreover, MGDebugger demonstrates remarkable debugging capabilities, particularly with DeepSeek-Coder-V2-Lite (16B) and Codestral (22B), where it achieves an accuracy of 94.5% on the HumanEval dataset, the highest score among all methods. This is especially impressive considering that MGDebugger operates in a zero-shot setting without task-specific retraining. This result illustrates the inherent debugging ability of larger LLMs with MGDebugger. Additionally, the method’s performance on MBPP, achieving an RSR of up to 41.1% with smaller models like CodeQwen1.5 (7B), further underscores its robustness. Statistical analysis using the Wilcoxon signed-rank test with 10 repetitions confirms that these improvements are significant (p ¡ 0.005). In general, these results validate MGDebugger as a highly effective and scalable debugging method for LLM-generated code.
Table 2. Results (Acc.) on HumanEval with advanced models.
Method Model
GPT-4o Claude 3.5 Sonnet LLaMA 3.1 70B
No-Debugging 90.2 89.6 79.3
Self-Debugging (Expl.) 93.3 (+3.1) 93.3 (+3.7) 86.6 (+7.3)
Reflexion 94.5 (+4.3) 93.9 (+4.3) 89.6 (+10.4)
MGDebugger 96.3 (+6.1) 95.7 (+6.1) 92.7 (+13.4)
To further validate the effectiveness and generalizability of MGDebugger across a wider spectrum of language models, we extended our evaluation to include recent advanced LLMs with strong natural language capabilities, including GPT-4o, Claude 3.5 Sonnet, and LLaMA 3.1 70B. We compare MGDebugger with the most competitive methods, Self-Debugging (Expl.) and Reflexion, on the HumanEval dataset.
The results in Table 2 demonstrate that MGDebugger continues to outperform existing debugging approaches even on state-of-the-art proprietary models like GPT-4o and Claude 3.5 Sonnet, which already possess strong zero-shot code generation capabilities. With these advanced models, MGDebugger achieves impressive accuracy improvements of +6.1% for both GPT-4o and Claude 3.5 Sonnet, pushing their overall accuracy to 96.3% and 95.7% respectively on HumanEval. This represents a substantial improvement over strong baselines like Self-Debugging and Reflexion. Notably, the results on LLaMA 3.1 70B demonstrate that MGDebugger provides even greater benefits for open-source models with enhanced natural language capabilities. The method yields a remarkable +13.4% improvement in accuracy, bringing LLaMA 3.1 70B’s performance to 92.7%.
These findings strengthen our previous conclusions about the adaptability of MGDebugger across different model architectures. We also reveal an interesting trend: while state-of-the-art models like GPT-4o and Claude 3.5 Sonnet start with higher baseline performance, MGDebugger still provides substantial improvements, suggesting that even the most advanced LLMs benefit significantly from structured debugging approaches. This highlights the potential for MGDebugger as a general enhancement method that can complement the evolving landscape of AI coding assistants.
3.3. Ablation Study (RQ2)
To understand the contribution of each component in MGDebugger and validate our design choices, we conduct an ablation study by systematically removing key components of our method: hierarchical code decomposition, LLM-simulated execution, and test case generation for subfunction debugging. Additionally, we explore the impact of replacing simulated execution with real execution traces collected with LDB (Zhong et al., 2024) to further validate our design choice. Each variant is evaluated on both the HumanEval and MBPP datasets using the DeepSeek-Coder-V2-Lite model.
Table 3. Results of ablation study.
@@ 3.3. Ablation Study (RQ2)
Acc. (%) RSR (%) Acc. (%) RSR (%)
MGDebugger 94.5 (+17.7) 76.3 80.0 (+12.8) 39.0
- w/o Hierarchical Debugging 89.0 (+12.2) 52.6 78.2 (+11.0) 33.5
- w/o Simulated Execution 90.2 (+13.4) 61.3 79.2 (+12.0) 36.6
- w/o Test Case Generation 90.9 (+14.1) 60.5 79.2 (+12.0) 36.6
- w/ Real Execution 92.7 (+15.9) 47.4 79.0 (+11.8) 36.0
@@
Table 4. Performance (RSR) on different bug categories in HumanEvalFix. The best and second-best scores are highlighted in bold and underline, respectively.
@@
Furthermore, we observe that replacing simulated execution with real execution leads to a noticeable drop in performance, particularly in repair success rate (from 76.3% to 47.4% on HumanEval and from 39.0% to 36.0% on MBPP). This suggests that real execution traces, which include all detailed variable changes and execution steps, may introduce noise that distracts the model during debugging. In contrast, simulated execution allows MGDebugger to focus on key variables and express debugging steps in a more reasoning-oriented manner, leading to better code understanding and more accurate repairs. Furthermore, simulated execution offers greater flexibility for extension to complex programming environments where execution environments may not be easily accessible (Ni et al., 2024). This makes MGDebugger more adaptable to a wider range of scenarios.
3.4. Debugging Different Types of Bugs (RQ3)
To assess the versatility and effectiveness of MGDebugger across various bug categories, we carry out experiments using the HumanEvalFix dataset, which is specifically designed to evaluate code debugging performance. The dataset involves six distinct bug categories: value misuse, missing logic, excess logic, operator misuse, variable misuse, and function misuse, allowing us to examine how effectively MGDebugger addresses different types of programming errors compared to existing methods.
(a) HumanEval dataset
(b) MBPP dataset
(c) HumanEvalFix dataset
Figure 3. Repair success rate of different methods when debugging code of different lengths.
(a) HumanEval dataset
(b) MBPP dataset
(c) HumanEvalFix dataset
Figure 4. Impact of debug attempts on the cumulative repair success rate of different methods.
Table 4 presents the RSRs across various bug categories with. We observe that MGDebugger consistently outperforms other methods with significantly higher overall accuracies. Specifically, MGDebugger achieves a remarkable repair success rate of 97.6% using DeepSeek-Coder, with 100% success rates in all bug categories except for value misuse. This is particularly notable given the complexity and diversity of the bugs in the dataset. Statistical analysis using the Wilcoxon signed-rank test with 10 repetitions confirms that these improvements are significant (p ¡ 0.005). This highlights the effectiveness of the hierarchical debugging strategy.
@@ 3.4. Debugging Different Types of Bugs (RQ3)
3.5. Performance Consistency (RQ4)
@@
We further assess the versatility of MGDebugger in debugging code of varying lengths (i.e., number of tokens), since code length often correlates with complexity and debugging challenges. We categorize code snippets into short, medium, and long groups based on the 1/3 and 2/3 quantiles of code length distribution, ensuring equal sample sizes across the three categories. We subsequently analyze the repair success rate scores obtained by MGDebugger and baselines when using DeepSeek-Coder as the backbone LLM.
@@
The results are presented in Figure 3. We observe that as the code length increases, most methods experience an obvious decrease in performance due to the increased complexity. Notably, MGDebugger consistently outperforms other methods across different code lengths on all three datasets. This performance advantage is particularly pronounced when debugging longer and more complex code snippets on HumanEval and HumanEvalFix, where MGDebugger maintains higher repair success rates compared to baseline methods.
Table 5. Number of bugs fixed in Defects4J.
Method Chart Closure Lang Math Mockito Time V1.2 V2.0
CodexRepair 9 30 22 29 6 3 99 31
GAMMA 11 24 16 25 3 3 82 45
ThinkRepair 11 31 19 27 6 4 98 47
ContrastRepair 12 32 19 30 8 2 103 40
Mulpor 13 26 19 23 7 4 92 59
ChatRepair 15 37 21 32 6 3 114 48
MGDebugger 13 40 27 40 9 5 129 64
@@
Another important factor for LLM-based debugging is the number of debugging attempts. Iterative debugging retries allows LLMs to refine their corrections over multiple passes, potentially leading to better outcomes (Yin et al., 2024; Zhong et al., 2024). We aim to assess MGDebugger’s ability to improve over successive iterations. Following Zhong et al. (2024), we vary the number of debugging attempts from 1 to 10 using DeepSeek-Coder on different dataset and collect the cumulative repair success rate (RSR) scores for each method. We clarify that a single iteration refers to one complete bottom-up traversal of the function decomposition tree, including test case generation and simulated execution for all subfunctions. Even if a function is decomposed into multiple subfunctions, the process is counted as a single iteration if it involves one full pass from the leaves to the root.
The results in Figure 4 show that MGDebugger achieves the highest cumulative RSR score among all methods. As shown in Figure 4(a), MGDebugger demonstrates steady improvement across iterations on HumanEval dataset, effectively helping the model correct its own generation errors and reaching a significantly higher success rate compared to baselines. For MBPP in Figure 4(b), MGDebugger shows consistent gains with each attempt, highlighting its ability to systematically identify and fix bugs in LLM-generated code. Figure 4(c) shows that on HumanEvalFix, MGDebugger achieves over 95% accuracy after 5 attempts and continues to improve gradually to reach 97.6%, though the improvement rate slows due to a few particularly challenging cases requiring specialized handling.
These results demonstrate MGDebugger’s strong potential in enhancing LLM-based code generation systems. By effectively debugging LLM-generated code through multiple iterations, MGDebugger not only improves the immediate code quality but also shows promise in strengthening the overall reliability of AI code generation systems. This iterative refinement capability is particularly valuable as it allows LLMs to learn from and correct their own mistakes, potentially leading to more robust and dependable automated programming solutions.
Figure 5. Bug fix Venn diagram in Defects4J V1.2.
3.6. Generalization to Real-World Software Defects (RQ5)
While our primary focus is to explore LLMs’ ability to fix LLM-generated code, we also evaluate MGDebugger on repairing real-world, human-written bugs. We leverage the widely used Defects4J dataset (Just et al., 2014), which contains real-world defects from open-source Java projects. We compare MGDebugger with state-of-the-art program repair approaches, including ChatRepair (Xia and Zhang, 2024), Mulpor (Lin et al., 2024), ContrastRepair (Kong et al., 2025), ThinkRepair (Yin et al., 2024), GAMMA (Zhang et al., 2024a), and CodexRepair (Xia et al., 2023). Following prior APR tools (Kong et al., 2025; Xia and Zhang, 2024; Lin et al., 2024; Yin et al., 2024), we use the 255 single-function bugs in Defects4J v1.2 and 82 single-line bugs in Defects4J v2.0 for evaluation and provide the perfect fault localization for each bug. To ensure a fair comparison, we use gpt-3.5-turbo-0301 as the LLM backbone and set the maximum number of generated patches to 200 following prior works (Kong et al., 2025; Xia and Zhang, 2024). For patch validation, we follow the established methodology from prior works (Kong et al., 2025; Xia and Zhang, 2024; Lin et al., 2024; Yin et al., 2024; Zhang et al., 2024a) and conduct manual validation by three experienced engineers who independently assess each patch. A ”correct” patch is defined as one that is semantically equivalent to the ground truth patch, while a ”plausible” patch passes all available test cases but may not fix the underlying bug. The engineers discuss inconsistencies and reach consensus through detailed examination of patch semantics and equivalence to ground truth solutions.
Table 5 presents the number of bugs fixed by MGDebugger and various state-of-the-art program repair approaches across different projects in Defects4J. MGDebugger demonstrates superior overall performance, successfully fixing 129 bugs in Defects4J v1.2 and 64 bugs in Defects4J v2.0. This represents an improvement of 15 more bugs (13.2%) over ChatRepair on Defects4J v1.2 and 5 more bugs (8.5%) over Mulpor on Defects4J v2.0. Notably, MGDebugger outperforms all other approaches on 5 out of 6 individual projects, with particularly strong absolute improvements.
Figure 5 illustrates the overlap of bugs fixed by MGDebugger and four strong baseline approaches (using ChatRepair results from the replication package of ContrastRepair (Kong et al., 2025) due to unavailable bug id details in the original paper (Xia and Zhang, 2024)). The diagram shows 36 bugs fixed by all methods, while MGDebugger uniquely fixes 25 distinct bugs, demonstrating its capability with complex defects. This complementarity with other approaches (e.g., Mulpor’s 26 uniquely fixed bugs) suggests potential for combining methods to achieve even higher repair rates, as different debugging strategies appear effective for various bug types and code contexts.
Overall, these results demonstrate that MGDebugger generalizes well beyond debugging LLM-generated code to fixing real-world software defects in production systems. The hierarchical decomposition and systematic debugging approach of MGDebugger proves effective for human-written code with complex project structures and dependencies. This generalizability is particularly important as it suggests that MGDebugger could serve as a unified debugging framework applicable to both AI-generated and human-written code.
Figure 6. Examples of code debugging by various methods on HumanEvalFix with DeepSeek-Coder.
4. Discussion
4.1. Case Study